### v6

@@ 4.3 Sketch of the Proof
(28) $\int_{\{G(v)\geq u\}}\dfrac{\|v-v^{*}\|_{2}}{\|\omega_{\alpha}\|_{L^{1}(\rho_{T_{\alpha,\epsilon}})}}e^{-\alpha\mathcal{E}(v)}\,d\rho_{T_{\alpha,\epsilon}}(v)\leq e^{\alpha(\delta)(\overline{\mathcal{E}}-\underline{\mathcal{E}})}\sqrt{2\mathcal{V}\big{(}\rho_{0}\big{)}}\cdot\dfrac{1}{\sqrt{u(\delta)}}\cdot\sqrt{\int G\,d\rho_{0}(v)}.$
@@
$\int_{\{G(v)\geq u\}}\dfrac{\|v-v^{*}\|_{2}}{\|\omega_{\alpha}\|_{L^{1}(\rho_{T_{\alpha,\epsilon}})}}e^{-\alpha\mathcal{E}(v)}\,d\rho_{T_{\alpha,\epsilon}}(v)$ $\leq e^{\alpha(\delta)(\overline{\mathcal{E}}-\underline{\mathcal{E}})}\sqrt{2\mathcal{V}\big{(}\rho_{0}\big{)}}\cdot\dfrac{1}{\sqrt{u(\delta)}}\cdot\sqrt{\int G\,d\rho_{0}(v)}$
@@
Thus we have proved that if all the conditions in Theorem 4.1 are satisfied, the desired decay can be achieved with the specified choices of $\alpha$ and $\epsilon$.
5 Numerical Experiments
@@ 5.1 Algorithm
$V^{j}_{k+1}=V^{j}_{k}-\lambda\gamma(V^{j}_{k}-v_{\alpha}(\hat{\rho}_{k}))-\left(\frac{\gamma}{\epsilon}\nabla G(V^{j}_{k})+\frac{\gamma}{\epsilon}\nabla^{2}G(V^{j}_{k})(V_{k+1}^{j}-V^{j}_{k})\right)-\sigma\sqrt{\gamma}(V^{j}_{k}-\bar{v}_{k})\odot z_{k},$
@@
(30) $V^{j}_{k+1}=V^{j}_{k}-\left[I+\frac{\gamma}{\epsilon}\nabla^{2}G(V^{j}_{k})\right]^{-1}\left(\lambda\gamma(V^{j}_{k}-v_{\alpha}(\hat{\rho}_{k}))+\frac{\gamma}{\epsilon}\sum_{i=1}^{m}g_{i}(V^{j}_{k})\nabla g_{i}(V^{j}_{k})+\sigma\sqrt{\gamma}(V^{j}_{k}-\bar{v}_{k})\odot z_{k}\right),$
@@
We propose an alternative algorithm when the dimensionality is high, where we introduce independent noise after the particles concentrate. This algorithm introduces additional noises to help the particles explore the landscape better, which is necessary when the dimension of the optimization problem is high. The complete algorithm is formulated as in Algorithm 2.
@@
$V^{j}\leftarrow V^{j}-\left[I+\frac{\gamma}{\epsilon}\sum_{i=1}^{m}\nabla^{2}\left[g^{2}_{i}(V^{j})\right]\right]^{-1}\left(\lambda\gamma(V^{j}-v_{\alpha}(\hat{\rho}))+\frac{\gamma}{\epsilon}\sum_{i=1}^{m}\nabla\left[g_{i}^{2}(V^{j})\right]+\sigma\sqrt{\gamma}(V^{j}-v_{\alpha}(\hat{\rho}))\odot z\right),$
@@
5: Output $)v_{\alpha}\left(\hat{\rho}\right),\mathcal{E}(v_{\alpha}\left(\hat{\rho}\right$))
@@
1:  while $|\mathcal{E}(v_{\alpha}\left(\hat{\rho}\right))-\mathcal{E}^{\star}|\geq\epsilon_{\text{indep}}$ do
@@
$V^{j}\leftarrow V^{j}-\left[I+\frac{\gamma}{\epsilon}\sum_{i=1}^{m}\nabla^{2}\left[g^{2}_{i}(V^{j})\right]\right]^{-1}\left(\lambda\gamma(V^{j}-v_{\alpha}(\hat{\rho}))+\frac{\gamma}{\epsilon}\sum_{i=1}^{m}\nabla\left[g_{i}^{2}(V^{j})\right]+\sigma\sqrt{\gamma}(V^{j}-v_{\alpha}(\hat{\rho}))\odot z\right),$
@@
5.2 Numerical examples
5.2.1 A simple example
We first test the algorithm on a $2$-dimensional example,
(31) $\min_{(v_{1},v_{2})\in\mathbb{R}^{2}}$ $v_{1}^{2}+v_{2}^{2}$
We test two difference types of constraints. The first case is an ellipse,
@@ 5.2.1 A simple example
The second case is a line,
@@
The exact minimizers are,
@@
We use Algorithm 1 with $N=50,\ \alpha=50,\ \epsilon=0.01,\ \lambda=1,\ \sigma=5,\ \gamma=0.1,\ \epsilon_{\text{stop}}=10^{-14},$ and the particles are initially set to follow a uniform distribution in the range of [3,3][-3,3] for both dimensions. We consider our search for the constrained minimizer successful if, when the algorithm finishes, $\|v_{\alpha}\left(\hat{\rho}\right)-v^{*}\|_{\infty}\leq 0.1$. The success rate and the average distance are shown in Table 1, where the average distance to $v^{*}$ in the table is measured using the following norm
@@
In Figure 4, we show the evolution of the objective function value $)\mathcal{E}(v_{\alpha}\left(\hat{\rho}\right$)), the constraint value $)g(v_{\alpha}\left(\hat{\rho}\right$)), and the distance $D(v_{\alpha}\left(\hat{\rho}\right),v^{*})$ over $100$ simulations. It is evident that the consensus point converges within 10 steps for all simulations.
In Figure 5, the evolution of all the particles and the consensus point are shown in time steps $k=0,5,50,100$. In all cases, after $5$ steps, most of the particles are driven to the constraints by the strong constraint term $\frac{1}{\epsilon}\nabla G(v)$ and stay there consistently. It is worth noting that in the case of the ellipse, not all particles converge around the consensus point. Some particles remain at the point $\tilde{v}$ where $\nabla g(\tilde{v})=0$ instead of satisfying $g(v)=0$. This happens when $G(v)=g^{2}(v)$ does not satisfy Assumption 4.2 (B1). However, it will not affect the convergence of the consensus point as long as the loss function value at $\tilde{v}$ is not significantly small compared to the constrained minimum. (See Remark 4.3 for more explanation.)
Table 1: The result of Algorithm 1 on (31) with constraints (32) or (33)
@@
(a) blue lines: $)\mathcal{E}(v_{\alpha}\left(\hat{\rho}\right$)), red lines: $)g(v_{\alpha}\left(\hat{\rho}\right$))
(b) $D(v_{\alpha}\left(\hat{\rho}\right),v^{*})$
(c) blue lines: $)\mathcal{E}(v_{\alpha}\left(\hat{\rho}\right$)), red lines: $)g(v_{\alpha}\left(\hat{\rho}\right$))
(d) $D(v_{\alpha}\left(\hat{\rho}\right),v^{*})$
Figure 4: The first line is the result for the optimization problem (31) with ellipse constraint (32), while the second line is for the line constraint (33). The left column is the evolution of the objective function value and constraint value, while the right column is the evolution of the distance between the consensus point and the exact minimizer, where the distance is defined in (34). The light lines are results from $100$ simulations, while the dark lines are the average values.
@@
Figure 5: The evolution of all the particles (in blue) and its consensus point (in red) in $2$-dimensional plane when solving for the constrained problem (31) with the ellipse constraint (32) (the first line) and the line constraint (33) (the second line). The constrained line is plotted in black, the black point is the global minimizer of the objective function, while the green point is the constrained minimizer $v^{*}$.
5.2.2 Ackley function
@@
where $b=1,A=20,a=0.1$, and $\hat{v}$ is the global minimum of the unconstrained problem. The above function in two-dimension is shown in Figure 1. Here we consider three different constraints,
(36) $\text{Case 1.}\quad\left\lVert v\right\rVert^{2}_{2}-1=0.$
(37) $\text{Case 2.}\quad\sum_{i=1}^{d-1}v_{i}^{2}-v_{d}=0.$
(38) $\text{Case 3.}\quad\sum_{i=1}^{d}v_{i}-1=0,\quad\displaystyle 2\sum_{i=1}^{d-1}v_{i}-\frac{1}{2}v_{d}-\frac{1}{2}=0.$
We set $\hat{v}=(0.4,\cdots,0.4)$, s.t. the unconstrained minimizer is not the same as the constrained minimizer. The constrained minimizers for the three-dimensional cases are
$\text{Case 1. }v^{*}=1/\sqrt{3}(1,1,1);\quad\text{Case 2. }v^{*}=(0.4283,0.4283,0.3669);\quad\text{Case 3. }v^{*}=(0.2,0.2,0.6).$
@@
$\text{Case 1. }v^{*}_{i}=1/\sqrt{20},\ 1\leq i\leq 20;\quad\text{Case 2. }v^{*}_{i}=0.3542,\ 1\leq i\leq 19,\ v^{*}_{20}=2.3839.$
@@
The evolution of the distance $D(v_{\alpha}\left(\hat{\rho}\right),v^{*})$ between the consensus point and the accurate solution is shown in Figure 6, where one can see that the consensus point converges to the true minimizer within $100$ steps. Besides, the success rate, averaged distance for the output consensus point $v^{*}$, and the averaged total steps are stated in Table 2. We consider the simulation to be successful if $\max_{k}|v_{\alpha}\left(\hat{\rho}\right)_{k}-v^{*}_{k}|\leq 0.1$, and the distance to $v^{*}$ is measured in $D(v_{\alpha}\left(\hat{\rho}\right),v^{*})$ and averaged over $100$ simulations. One can see that except for the $20$-dimensional case 2, the algorithm can find the exact minimizer within $400$ steps with $100\%$ success rate. Even for the $20$-dimensional case 2, although the success rate is a bit less than $100\%$, the average distance to $v^{*}$ is less than $0.05$, which means that they are all relatively close to the exact minimizer $v^{*}$.
The reason for the nonsmoothness in the later stage of the average line is due to the limited number of samples for the larger steps. In most simulations, the algorithm typically concludes its iterations around the average total steps in the table. As it is hard to find the exact minimizer for $20$-dimensional Ackley function with the constraints (38), so we only plot the result for case 3 in $3$-dimension.
Table 2: The result of Algorithm 1 on $3$-dimensional Ackley function and Algorithm 2 for $20$-dimensional Ackley function.
@@
case 1 d=3 $100\%$ $8\times 10^{-3}$ 295
@@
(b) $d=3$ case 2
(c) $d=3$ case 3
(d) $d=20$ case 1
@@
Figure 6: The evolution of the distance $D(v_{\alpha}\left(\hat{\rho}\right),v^{*})$ between the consensus point and the exact minimizer. The objective function is the Ackley fuction (35) and the constraints are (36) - case3. The light lines are the results from $100$ simulations, while the dark lines are the average values.
To maintain clarity, we present only the two numerical examples above in the main text. An additional example, focusing on the Thompson problem, is provided in Appendix J to further illustrate the scalability and effectiveness of the proposed algorithm in high-dimensional settings involving up to thousands of variables.
6 Conclusions
In this paper, we propose a new CBO-based method for solving constrained non-convex minimization problem with equality constraints and potentially non-differentiable loss functions. Specifically, we augment the original CBO framework with a new forcing term designed to guide particles toward the constraint set. On the theoretical side, we conduct a rigorous analysis of the mean-field limit for the proposed model (6), deriving the corresponding macroscopic model (7) and establishing well-posedness results for both the microscopic and macroscopic models. To demonstrate the convergence of the method, we study the long-time behavior of the macroscopic model (7) through an analysis of the associated Fokker-Planck equation (8). Our results establish that, under Assumption 4.2 and with a proper choice of parameters, particles converge to the constrained minimizer $v^{*}$ with arbitrary closeness. Notably, Assumption 4.2 (C) fits well with the basic nature of the algorithm, while Assumption 4.2 (B) serves as a technical requirement needed by our proof technique, which might be relaxed further with an alternative proof technique, as suggested by the performance exhibited in numerical experiments where Assumption 4.2 (B) may not be strictly satisfied. On the practical side, we proposed a stable algorithm based on the continuous-in-time model. In Section 5, the algorithm’s performance is illustrated through a series of experiments, including challenging high-dimensional problems.

### v8

@@ 4.3. Sketch of the Proof
(28) $\int_{\{G(v)\geq u\}}\dfrac{\|v-v^{*}\|_{2}}{\|\omega_{\alpha}\|_{L^{1}(\rho_{T_{\alpha,\epsilon}})}}e^{-\alpha\mathcal{E}(v)}\,d\rho_{T_{\alpha,\epsilon}}(v)$
$\leq e^{\alpha(\delta)(\overline{\mathcal{E}}-\underline{\mathcal{E}})}\sqrt{2\mathcal{V}\big(\rho_{0}\big)}\cdot\dfrac{1}{\sqrt{u(\delta)}}\cdot\sqrt{\int G\,d\rho_{0}(v)}.$
@@
$\int_{\{G(v)\geq u\}}\dfrac{\|v-v^{*}\|_{2}}{\|\omega_{\alpha}\|_{L^{1}(\rho_{T_{\alpha,\epsilon}})}}e^{-\alpha\mathcal{E}(v)}\,d\rho_{T_{\alpha,\epsilon}}(v)$
$\leq e^{\alpha(\delta)(\overline{\mathcal{E}}-\underline{\mathcal{E}})}\sqrt{2\mathcal{V}\big(\rho_{0}\big)}\cdot\dfrac{1}{\sqrt{u(\delta)}}\cdot\sqrt{\int G\,d\rho_{0}(v)}$
@@
Thus we have proved that if all the conditions in Theorem 4.1 are satisfied, the desired decay can be achieved with the specified choices of $\alpha$ and $\epsilon$. ∎
5. Numerical Experiments
@@ 5.1. Algorithm
$V^{j}_{k+1}=$ $V^{j}_{k}-\lambda\gamma(V^{j}_{k}-v_{\alpha}(\hat{\rho}_{k}))$
$-\left(\frac{\gamma}{\epsilon}\nabla G(V^{j}_{k})+\frac{\gamma}{\epsilon}\nabla^{2}G(V^{j}_{k})(V_{k+1}^{j}-V^{j}_{k})\right)-\sigma\sqrt{\gamma}(V^{j}_{k}-\bar{v}_{k})\odot z_{k},$
@@
(30) $V^{j}_{k+1}=V^{j}_{k}-\left[I+\frac{\gamma}{\epsilon}\nabla^{2}G(V^{j}_{k})\right]^{-1}\Bigg($ $\lambda\gamma(V^{j}_{k}-v_{\alpha}(\hat{\rho}_{k}))$
$+\frac{\gamma}{\epsilon}\sum_{i=1}^{m}g_{i}(V^{j}_{k})\nabla g_{i}(V^{j}_{k})+\sigma\sqrt{\gamma}(V^{j}_{k}-\bar{v}_{k})\odot z_{k}\Bigg),$
@@
We also propose an alternative algorithm when the dimensionality is high, where we introduce independent noise after the particles concentrate. This version introduces additional noises to help the particles explore the landscape better, which is necessary when the dimension of the optimization problem is high. The complete algorithm is formulated as in Algorithm 2.
@@
$\displaystyle V^{j}\leftarrow V^{j}-\left[I+\frac{\gamma}{\epsilon}\sum_{i=1}^{m}\nabla^{2}\left[g^{2}_{i}(V^{j})\right]\right]^{-1}\Bigg($ $\lambda\gamma(V^{j}-v_{\alpha}(\hat{\rho}))+\frac{\gamma}{\epsilon}\sum_{i=1}^{m}\nabla\left[g_{i}^{2}(V^{j})\right]$
$+\sigma\sqrt{\gamma}(V^{j}-v_{\alpha}(\hat{\rho}))\odot z\Bigg),$
@@
5:Output $)v_{\alpha}\big(\hat{\rho}\big),\mathcal{E}(v_{\alpha}\big(\hat{\rho}\big$))
@@
1:while $|\mathcal{E}(v_{\alpha}\big(\hat{\rho}\big))-\mathcal{E}^{\star}|\geq\epsilon_{\text{indep}}$ do
@@
$\displaystyle V^{j}\leftarrow V^{j}-\left[I+\frac{\gamma}{\epsilon}\sum_{i=1}^{m}\nabla^{2}\left[g^{2}_{i}(V^{j})\right]\right]^{-1}\Bigg($ $\lambda\gamma(V^{j}-v_{\alpha}(\hat{\rho}))+\frac{\gamma}{\epsilon}\sum_{i=1}^{m}\nabla\left[g_{i}^{2}(V^{j})\right]$
$+\sigma\sqrt{\gamma}(V^{j}-v_{\alpha}(\hat{\rho}))\odot z\Bigg),$
@@
5.2. Numerical examples
5.2.1. A simple example
We first test the proposed algorithm on a simple objective function,
(31) $\min_{v\in\mathbb{R}^{d}}$ $\|v\|_{2}^{2}.$
We consider three types of constraints. The first is a line-segment constraint in $\mathbb{R}^{3}$ ($d=3$), with
(32) $g(v)=\mathrm{dist}(v,L),$
where $\mathrm{dist}(\cdot,L)$ denotes the Euclidean distance to the line segment $L$ with endpoints $(1.6,0.2,0.4)$ and (0.3,0.7,0.5)(-0.3,-0.7,0.5). The second case is an ellipse, with
@@ 5.2.1. A simple example
The third case is a line, with
@@
The exact minimizers of those three cases are,
${\text{Line-segment: }v^{*}=(0.2361,-0.446,0.4718)};$
@@
We use Algorithm 1 with $N=50,\ \alpha=50,\ \epsilon=0.01,\ \lambda=1,\ \sigma=5,\ \gamma=0.1,\ \epsilon_{\text{stop}}=10^{-14},$ and the particles are initially set to follow a uniform distribution in $3,3]^{d}$. We consider our search for the constrained minimizer successful if, when the algorithm finishes, $\|v_{\alpha}\big(\hat{\rho}\big)-v^{*}\|_{\infty}\leq 0.1$. The success rate and the average distance are shown in Table 1, where the average distance to $v^{*}$ in the table is measured using the following norm
@@
In Figure 4, we show the evolution of the objective function value $)\mathcal{E}(v_{\alpha}\big(\hat{\rho}\big$)), the constraint value $)g(v_{\alpha}\big(\hat{\rho}\big$)), and the distance $D(v_{\alpha}\big(\hat{\rho}\big),v^{*})$ over $100$ simulations. It is evident that the consensus point stabilizes within 10 steps for all simulations.
In Figure 5, the evolution of all the particles and the consensus point are shown in time steps $k=0,5,50,100$. In all cases, after $5$ steps, most of the particles are driven to the constraints by the strong constraint term $\frac{1}{\epsilon}\nabla G(v)$ and stay there consistently. It is worth noting that, in contrast to the line-segment case, which satisfies Assumption 2 (B) (guaranteed by Lemma 4.2), in the ellipse case not all particles converge to the consensus point. Instead, some particles remain trapped at a point $\tilde{v}$ satisfying $\nabla g(\tilde{v})=0$, rather than reaching the feasible set $\{v:g(v)=0\}$. This behavior occurs when $G(v)=g^{2}(v)$ fails to satisfy Assumption 2 (B), in particular condition (B1). Nevertheless, it does not affect the convergence of the consensus point as long as the loss value at $\tilde{v}$ is not too small compared to the constrained minimum (see Remark 3).
Table 1. The result of Algorithm 1 on (31) with constraints (32), (33), and (34)
@@
(a) blue lines: $)\mathcal{E}(v_{\alpha}\big(\hat{\rho}\big$)), red lines: $)g(v_{\alpha}\big(\hat{\rho}\big$))
(b) $D(v_{\alpha}\big(\hat{\rho}\big),v^{*})$
(c) blue lines: $)\mathcal{E}(v_{\alpha}\big(\hat{\rho}\big$)), red lines: $)g(v_{\alpha}\big(\hat{\rho}\big$))
(d) $D(v_{\alpha}\big(\hat{\rho}\big),v^{*})$
(e) blue lines: $)\mathcal{E}(v_{\alpha}\big(\hat{\rho}\big$)), red lines: $)g(v_{\alpha}\big(\hat{\rho}\big$))
(f) $D(v_{\alpha}\big(\hat{\rho}\big),v^{*})$
Figure 4. The first row is the result for the optimization problem (31) with line-segment constraint (32), the second row is the result for the ellipse constraint (33), and the third row is for the line constraint (34). The left column is the evolution of the objective function value and constraint value, while the right column is the evolution of the distance between the consensus point and the exact minimizer, where the distance is defined in (35). The light lines are results from $100$ simulations, while the dark lines are the average values.
@@
(i) k=0
(j) k=5
(k) k=50
(l) k=100
Figure 5. Evolution of all particles (blue) and the corresponding consensus point (red) for the constrained problems: (32) with the line-segment constraint (first two rows), (31) with the ellipse constraint (33) (third row), and (31) with the line constraint (34) (fourth row). The constraint set is shown in black. The black marker denotes the unconstrained global minimizer of the objective function, and the green marker denotes the constrained minimizer $v^{*}$.
5.2.2. Ackley function
@@
where $b=1,A=20,a=0.1$, and $\hat{v}$ is the global minimum of the unconstrained problem. The above function in two-dimension is shown in Figure 1. Here we consider five different constraints,
[…]
(37) $(0.2,0.5,0.7)$ and $(0.5,0.2,0.5)$
[…]
(38) $\{v\mid v_{1}^{2}+\cdots+v_{18}^{2}\leq 0.25,\ v_{19}=v_{20}=0\}$
(39) $\text{Case 3.}\quad\|v\|_{2}^{2}-1=0\ \text{in $\mathbb{R}^{3}$ and in $\mathbb{R}^{20}$}$
(40) $\text{Case 4.}\quad\sum_{i=1}^{d-1}v_{i}^{2}-v_{d}=0\ \text{in $\mathbb{R}^{3}$ and in $\mathbb{R}^{20}$}$
(41) $\text{Case 5.}\quad g(v)=\sum_{i=1}^{d}v_{i}-1=0,\quad\displaystyle 2\sum_{i=1}^{d-1}v_{i}-\frac{1}{2}v_{d}-\frac{1}{2}=0\ \text{in $\mathbb{R}^{3}$}$
By by Lemma 4.2, Case 1 and Case 2 satisfy Assumption 2 (B). Here, we set $\hat{v}=(0.4,\cdots,0.4)$ so that the unconstrained minimizer is not the same as the constrained minimizer. The constrained minimizers for the three-dimensional cases are
${\text{Case 1. }v^{*}=(0.4003,0.2997;0.5665);}\quad\text{Case 3. }v^{*}=1/\sqrt{3}(1,1,1);$
$\text{Case 4. }v^{*}=(0.4283,0.4283,0.3669);\quad\text{Case 5. }v^{*}=(0.2,0.2,0.6).$
@@
${\text{Case 2. }v_{i}^{*}=0.1179,\ 1\leq i\leq 18,\ v_{19}^{*}=v_{20}^{*}=0};$
$\text{Case 3. }v_{i}^{*}=1/\sqrt{20},\ 1\leq i\leq 20;$
$\text{Case 4. }v_{i}^{*}=0.3542,\ 1\leq i\leq 19,\ v_{20}^{*}=2.3839.$
@@
${\text{Case 2.}\ \epsilon_{\min}=0.05,\ \sigma_{\mathrm{indep}}=0.3;}$
@@
The evolution of the distance $D(v_{\alpha}\big(\hat{\rho}\big),v^{*})$ between the consensus point and the accurate solution is shown in Figure 6, where one can see that the consensus point converges to the true minimizer within $100$ steps. Besides, the success rate, averaged distance for the output consensus point $v^{*}$, and the averaged total steps are reported in Table 2. We consider the simulation to be successful if $\max_{k}|v_{\alpha}\big(\hat{\rho}\big)_{k}-v^{*}_{k}|\leq 0.1$, and the distance to $v^{*}$ is measured in $D(v_{\alpha}\big(\hat{\rho}\big),v^{*})$ and averaged over $100$ simulations. One can see that, except for the $20$-dimensional Cases 2 and 4, the algorithm finds the exact minimizer with a $100\%$ success rate. Moreover, except for the $3$-dimensional Case 1 and the $20$-dimensional Case 4, the algorithm achieves high accuracy within $400$ steps. Even for the $20$-dimensional Cases 2 and 4, although the success rate is slightly below $100\%$, the average distance to $v^{*}$ is below $0.05$, indicating that the iterates remain relatively close to the exact minimizer $v^{*}$.
The reason for the non-smoothness in the later stage of the average line is due to the limited number of samples for the larger steps. In most simulations, the algorithm typically concludes its iterations around the average total steps in the table. As it is difficult to obtain the exact constrained minimizer of the $20$-dimensional Ackley function under the constraints in (41), we only consider Case 5 in $\mathbb{R}^{3}$. Moreover, for the $20$-dimensional Case 2, $v^{*}$ is a reference solution computed by first projecting the unconstrained global minimizer onto $S$ and then using the MATLAB Optimization Toolbox to refine the solution, since the exact constrained minimizer is hard to obtain.
Table 2. The result of Algorithm 1 on 33-dimensional Ackley function and Algorithm 2 for 2020-dimensional Ackley function.
@@
case 1 d=3 $100\%$ $3.8\times 10^{-3}$ 1000
case 2 d=20 $98\%$ $2.4\times 10^{-2}$ 158
case 3 d=3 $100\%$ $8\times 10^{-3}$ 295
@@
(b) $d=3$ case 3
(c) $d=3$ case 4
(d) $d=3$ case 5
@@
(f) $d=20$ case 3
(g) $d=20$ case 4
Figure 6. The evolution of the distance $D(v_{\alpha}\big(\hat{\rho}\big),v^{*})$ between the consensus point and the exact minimizer. The objective function is the Ackley fuction (36) and the constraints are (37) - (41). The light lines are the results from $100$ simulations, while the dark lines are the average values.
5.2.3. Thomson’s Problem
We use Algorithm 2 with
[…]
and all the particle initially follow $V^{j}\sim\text{Unif}[-1,1]^{3k}$.
[…]
(42) $\text{relative error}=\frac{|\mathcal{E}(v_{\alpha}\big(\hat{\rho}\big))-\mathcal{E}(v^{*})|}{\mathcal{E}(v^{*})}$
[…]
Table 3. The result of Algorithm 2 on Thomson problem.
success rate relative error constraints value total steps
$k=2,(d=6)$ $100\%$ $4.4\times 10^{-3}$ $3.8\times 10^{-11}$ $382$
$k=3,(d=9)$ $100\%$ $9.9\times 10^{-3}$ $1.4\times 10^{-10}$ $407$
$k=8,(d=24)$ $100\%$ $1.78\times 10^{-2}$ $2.3\times 10^{-10}$ $567$
$k=15,(d=45)$ $100\%$ $1.57\times 10^{-2}$ $3.4\times 10^{-10}$ $895$
$k=56,(d=168)$ $97\%$ $1.44\times 10^{-2}$ $2.91\times 10^{-6}$ $1610$
$k=470,(d=1410)$ $93\%$ $1.95\times 10^{-2}$ $4.03\times 10^{-6}$ $1960$
(a) k=2
(b) k=3
(c) k=8
(d) k=15
(e) k=56
(f) k=470
Figure 7. Thomson Problem: the decay of the relative error over $100$ simulation and its mean.
6. Conclusions
In this paper, we propose a new CBO-based method for solving constrained non-convex minimization problem with equality constraints and potentially non-differentiable loss functions. Specifically, we augment the original CBO framework with a new forcing term designed to guide particles toward the constraint set. On the theoretical side, we conduct a rigorous analysis of the mean-field limit for the proposed model (6), deriving the corresponding macroscopic model (7) and establishing well-posedness results for both the microscopic and macroscopic models. To demonstrate the convergence of the method, we study the long-time behavior of the macroscopic model (7) through an analysis of the associated Fokker-Planck equation (8). Our results establish that, under Assumption 2 and with a proper choice of parameters, particles converge to the constrained minimizer $v^{*}$ with arbitrary closeness. Notably, Assumption 2 (C) fits well with the basic nature of the algorithm, while Assumption 2 (B) serves as a technical requirement needed by our proof technique, which might be relaxed further with an alternative proof technique, as suggested by the performance exhibited in numerical experiments where Assumption 2 (B) may not be strictly satisfied. On the practical side, we proposed a stable algorithm based on the continuous-in-time model. In Section 5, the algorithm’s performance is illustrated through a series of experiments, including challenging high-dimensional problems.